[figure: 9.7 cm]

Ferromagnetic Monolayers: A fresh Look at Fundamentals

Klaus Baberschke^{a} aInstitut für Experimentalphysik, Freie Universität Berlin, Arnimallee 14, D-14195 Berlin-Dahlem, Germany

Ultrathin ferromagnetic films of Fe, Co, Ni, etc. allow manipulations of the magnetism which are hardly accessible in the bulk : (i) Due to special growth mechanisms crystallographic structures can be prepared which are impossible in bulk magnetism, e.g. tetragonal Ni single crystals. (ii) As a consequence of the finite size effect one can shift the phase transition temperature T_{C} to almost any value between zero and the bulk Curie-temperature. (i) Changes of the c/a-ratio by a few percent or ≈ 0.05 Å of the nearest neighbor distance will change the magnetic anisotropy energy by 10^{2} −10^{3}! This dramatic change is caused by the reactivation of orbital magnetism, being not anymore quenched like in bulk cubic ferromagnets. Experimental examples will be given which agree perfectly with ab-initio theory if the input parameter for ab-initio calculations are the real structures and not idealized fcc-, bcc-lattices. (ii) The thickness dependence of T_{C} is used to study the magnetic phase transition and the interlayer exchange coupling in a trilayer of Co/Cu/Ni. Separate T_{C}^{Ni} and T_{C}^{Co} for a coupled trilayer are observed and the effect of enhanced spin fluctuations in two dimensions will be discussed. For such an exchange-coupled trilayer we find a very strong temperature dependence of J_{inter}(T). Finally, we will discuss the Curie-Weiss-law for thin Fe films and its critical exponent γ. The work was supported by DFG, Sfb 290 and BMBF(05KS1 KEB4)

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : invited

Corresponding author : Klaus Baberschke

Address for correspondence : Institut fur Experimentalphysik, Freie Universitat Berlin, Arnimallee 14, D-14195 Berlin-Dahlem, Germany

Email address : bab@physik.fu-berlin.de